Project

General

Profile

10-August-2015

Attending
Satish, Paola, Joseph, David, Craig, Paul, Chris, Qiulan

News

We will be having a production workshop Aug 24-25.
We are working to incorporate MR Brem into the standard production chain.
Additional samples coming down the line:
  • Modern sim rock files (2M events)
  • ND MC for top-up ND data
  • NuE only ND files
  • Nu on electron ND MC
  • Flat energy flux ND MC

Satish will be on vacation Wednesday evening through Sunday evening. He expects to be unreachable for most of that time. There is a small chance this might extend though Monday.

Software Tags

David will cut a new tag tomorrow. Any new software commits for this tag should go in by noon tomorrow.

ND Additional First Analysis MC Generation (Paul)

Last week found out jobs not using genieRW. Over the weekend he submitted jobs using the version of the overlay script in SVN. However, jobs are still failing for reasons that Paul has not yet completely resolved. He may be sending emails asking for help soon.

Reco

Joseph has mostly finished reconstruction of additional ND "top-up" data. A small subset has not yet been finished because of a typo in the constraints defining the ending run-number. Even that is now processed to within a few files. He will create and send around new datasets. Joseph once again emphasized the need to rationalize our epoch naming scheme and to stick to it.

LEM

Chris has been unable to LEM the RHC Monte Carlo because of the missing diblock mask crash. The exception for ideal conditions MC has been backported to S15-05-07a, but it was evidently not published properly. David will investigate, but in the mean-time, this is stalled. Chris is caught up with the LEMing of the available ND top-up data, and will be finishing it up once he has inputs from Joseph.

CAF/Mix

Bruno is on vacation until Aug 12, so there is no update from him for the ND top-up.

raw2root keepup (Paola)

This is running without any major issues. One file crashed with error code 65. Poala will update the keepup ticket with more details

calibration keepup

There were some transient errors last week, related to failed database updates. These have been addressed, and calibration is now running smoothly again.

Reco keepup (Vito)

This is now complete. All the leftover files since April 1st have been reprocessed.

Files up through April 11 - June 20 were processed with S5-05-07a. Of these a total of 216 FD files failedwith known BPF crashes.
Files since June 20 have been processed with S15-07-30. These have all been successfully processed up through the start of the shutdown. No errors have been observed.

ND CRY calibration (Qiulan)

This is now complete. Some files didn't get through initially because FTS error, but retries successfully cleared those.

AOB:

Chris observes some missing NUMI files in watchdog plots. He will contact the watchdogs list.

Chris is working on script to merge subruns for summing caf's. The possibility of merging ART files was discussed, but we have no good solution for preventing file sizes from getting too big.